### v5

# A Convex Hull Cheapest Insertion Heuristic for the Non-Euclidean TSP

Abstract
The convex hull cheapest insertion heuristic produces good solutions to the Euclidean Traveling Salesperson Problem, but it has never been extended to the non-Euclidean problem. This paper uses multidimensional scaling to first project the points from a non-Euclidean space into a Euclidean space, enabling the generation of a convex hull that initializes the algorithm. To evaluate the proposed algorithm, non-Euclidean spaces are created by adding separators to the TSPLIB data-set, or by using the $\mathcal{L}_{1}$ norm as a metric.
keywords:
Traveling salesman problems, Vehicle routing and navigation, Control and Scheduling, Algorithms, Logistics, Supply Chains
1 Introduction
The Traveling Salesperson Problem (TSP) involves finding the shortest possible tour that visits a set of locations exactly once before returning to the starting location [1]. In contrast to TSPs in Euclidean spaces, the non-Euclidean TSP includes environmental obstacles, or a cost function that is not simply the pairwise Euclidean distance between locations [2]. For example, if the cost to be minimized is the time spent traveling between locations in a city, the optimal tour is dependent on the available road infrastructure that causes path deviations from the straight-line path between locations [3]. In case of a rectangular grid of streets, the $\mathcal{L}_{1}$ norm or the Manhattan norm is often used instead of the Euclidean distance [4, 5].
When the size of the problem increases, i.e., for a large number of locations, exact methods for computing the optimal solution to TSPs are intractable due to their $\mathcal{NP}$-hard nature [6]. To accelerate exact methods by providing a solution as a warm start [7], or when TSP solutions have to be found in real time, heuristic algorithms are used to rapidly find reasonably good solutions [8]. Among heuristics developed for Euclidean problems, the Convex Hull Cheapest Insertion (CHCI) heuristic has been shown to produce superior solutions to greedy heuristics [9, 10]. The CHCI heuristic is initiated by a subtour of the convex hull of locations, and its interior points are then progressively incorporated into the subtour in increasing order of insertion cost, until the complete tour is obtained. The initiation of the candidate subtour with the convex hull of the TSP points is advantageous because points on the boundary of the convex hull are visited in the same cyclic order as they appear in the optimal Euclidean TSP tour [11, 12].
The initialization of the CHCI heuristic by the convex hull of points has prevented the utilization of this promising heuristic in non-Euclidean TSPs. In general, it is noted that effective heuristics for the non-Euclidean TSP have been neglected in literature, and the simple Nearest Neighbor (NN) or Nearest Insertion (NI) heuristics are commonly used instead [13]. Another approach is to either neglect obstacles, or to replace the non-Euclidean cost function with Euclidean distances [14]. However, this approximation is not always acceptable, especially when the true costs deviate significantly from the Euclidean distance [15]. In this context, using methods developed for Euclidean TSPs results in sub-optimal solutions [13].
The contribution of this paper is the extension of the Euclidean CHCI algorithm to the non-Euclidean TSP, motivated by the expected reduction in tour cost when compared to the NN and NI heuristics. This is achieved by first applying multidimensional scaling (MDS) to find the set of points in a projected Euclidean space, whose pairwise distances approximate the non-Euclidean pairwise cost [16, 17]. The points on the boundary of their convex hull are used to initiate a sub-tour onto which the remaining points are added in increasing order of their true non-Euclidean insertion cost ratios. While MDS has recently been applied to the non-Euclidean TSP for tour length estimation, local clustering, and to understand human cognition [18, 19, 20], the use of MDS to initiate the CHCI heuristic for the non-Euclidean TSP is a novel approach.
After defining the proposed algorithm and its competing benchmark algorithms, this paper reports the outcomes of extensive computational experiments on TSPLIB benchmark instances that are modified by adding separators or using the $\mathcal{L}_{1}$ norm to set up non-Euclidean test cases [21]. This novel heuristic is demonstrated to outperform the NN and NI heuristic in 88% and 99% of the cases studied, respectively. When compared with metaheuristic algorithms, the proposed heuristic’s tour costs are lower than the solutions found by the genetic algorithm and ant colony optimization algorithm in 87% and 95% of the instances respectively.
2 Non-Euclidean TSP
Consider a complete directed graph represented by $\mathcal{G}:=(V,A)$ where $V$ is the set of locations or nodes, and the directed arc set $A:=\{(v_{i},v_{j})|v_{i},v_{j}\in V,\forall i\neq j\}$ which connects every ordered pair of distinct nodes in $V$. A cost function $c:A\rightarrow\mathbb{R}^{+}$ defines the metric to be minimized, and each arc $(v_{i},v_{j})\in A$ is associated with a defined cost $c(v_{i},v_{j})\in\mathbb{R}^{+}$. The objective of the TSP is to find the minimum cost sequence of arcs on the graph $\mathcal{G}$ that visits every node in $V$ exactly once and returns to the starting node. The resulting sequence is called a minimum cost Hamiltonian tour, and if $|V|=n$, it can be expressed as a sequence $T=(v_{1},v_{2},...,v_{n},v_{1})$ whose tour cost is the sum of costs of constituting arcs, given by $J=\textstyle\sum_{r=1}^{n}c(v_{r},v_{r+1})$. A constraint $v_{n+1}=v_{1}$ enforces that the starting and ending node are the same location.
When the cost function $c$ defines a non-Euclidean metric for the relation between every pair of nodes, finding a minimum cost Hamiltonian cycle is called the non-Euclidean TSP. Let the non-Euclidean arc costs be captured by a cost matrix $C\in\mathbb{R}^{n\times n}$ whose entries are defined as $C_{ij}=c(v_{i},v_{j})~{}\forall(v_{i},v_{j})\in A$.
3 Adapted Convex Hull Cheapest Insertion Algorithm
The proposed algorithm is referred to as the adapted CHCI (ACHCI) algorithm. It first uses MDS to find a set of points in 2D space whose pairwise Euclidean distance approximates the non-Euclidean arc cost function $c$. This is initiated by assigning one of the TSP points to be the origin in a new Euclidean space, and then finding the coordinates of the remaining $n-1$ points such that their pairwise Euclidean distances are exactly their non-Euclidean costs. Assuming a full rank cost matrix $C$, these points are first obtained in an $n-1$ dimension Euclidean space, after which they are projected to a 2D space.
Let the Euclidean coordinate equivalent of the node $v_{i}\in V$ be represented by $x_{i}\in\mathbb{R}^{n-1}$. In this paper, the notation $x_{i}\leftrightarrow v_{i}$ is used to indicate this mapping. For some $x_{j},x_{k}\in\mathbb{R}^{n-1}$, the relative position vectors of $x_{i}$ and $x_{j}$ with respect to $x_{k}$ are $(x_{i}-x_{k})$ and $(x_{j}-x_{k})$ respectively. If the Euclidean distance between these points is identical to their respective non-Euclidean costs as defined in cost matrix $C$, it can be verified that their inner product satsifies
$\langle x_{i}-x_{k},x_{j}-x_{k}\rangle=(C_{ik}^{2}+C_{kj}^{2}-C_{ij}^{2})/2$ (1)
This relation between the desired Euclidean coordinates and the cost matrix $C$ is leveraged to calculate the coordinates of each point relative to the origin in the $n-1$ dimensional Euclidean space. Without loss of generality, pick position vector $x_{1}\leftrightarrow v_{1}$ as the origin and define $\bar{x}_{i}:=x_{i}-x_{1}$ as the relative position vector of $x_{i}$ with respect to $x_{1}$. Let the ordered collection of all relative position vectors form a column matrix $\bar{X}\in\mathbb{R}^{(n-1)\times(n-1)}:=[\bar{x_{2}},\bar{x_{3}},...,\bar{x_{n}}]$.
Define the Gramian matrix $M\in\mathbb{R}^{(n-1)\times(n-1)}$ by $M_{ij}:=\langle\bar{x_{i}},\bar{x_{j}}\rangle$. Then by definition,
$M$ $=\bar{X}^{\top}\bar{X}$ (2)
Figure 1: Using multi-dimensional scaling to find the Euclidean 2D approximation of a non-Euclidean instance with 25 points
Each entry of matrix $M$ can be calculated using the cost matrix $C$ and Eq. (1). The eigenvalue decomposition of $M$ results in eigenvector matrix $Q\in\mathbb{R}^{(n-1)\times(n-1)}$ and diagonal eigenvalue matrix $\Lambda\in\mathbb{R}^{(n-1)\times(n-1)}$ as shown in Eq. (3). Because Gram matrices are positive semi-definite, the eigenvalues of $M$ are necessarily non-negative. If $\Sigma$ is the diagonal singular value matrix so that $\Lambda=\Sigma^{\top}\Sigma$, then,
$M=Q\Lambda Q^{\top}=Q\Sigma^{\top}\Sigma Q^{\top}$ $=(\Sigma Q^{\top})^{\top}(\Sigma Q^{\top})$ (3)
Using Eq. (2) and (3), it is clear that $\bar{X}=\Sigma Q^{\top}$ defines a set of point coordinates that have Euclidean distances that are identical to the defined non Euclidean costs. However, they are in an $n-1$ dimensional space, where, for $n>2$, obtaining the convex hull initiated TSP subtour is challenging. For this reason, principal component analysis is applied to obtain the set of two-dimensional point coordinates with pairwise distances that best approximate the non-Euclidean cost function. The two largest eigenvalues $\lambda_{max_{1}},\lambda_{max_{2}}$ are chosen to form $\widetilde{\Sigma}\in\mathbb{R}^{2\times 2}:=$ diag$(\sqrt{\lambda_{max_{1}}},\sqrt{\lambda_{max_{2}}})$. Let $\widetilde{Q}\in\mathbb{R}^{(n-1)\times 2}$ contain their respective eigenvectors.
The approximated 2D coordinates are then obtained as the columns of $\widetilde{X}\in\mathbb{R}^{2\times(n-1)}$:
$\widetilde{X}=\widetilde{\Sigma}\widetilde{Q}^{T}$
This procedure is illustrated for a set of 25 enumerated points that are separated by impassable line segments, creating a non-Euclidean point cloud as shown in Fig. LABEL:Im1a:_PointCloud. The 2D Euclidean approximate coordinates that define matrix $\widetilde{X}$ are shown in Fig. LABEL:Im1b:_MDSAppx where the pairwise Euclidean distances approximate the pairwise shortest true paths of the original non-Euclidean point cloud. Also shown is the convex hull of these 2D points that initiates the subtour of the ACHCI heuristic.
Figure 2: ACHCI subtour initiation and complete tour for a non-Euclidean TSP instance with 25 points
The ACHCI algorithm is summarized as:
  1. 1.
    Use MDS to obtain 2D Euclidean approximate coordinates that define matrix $\widetilde{X}$
  2. 2.
    Initiate subtour as the convex hull nodes of $\widetilde{X}$. Next, discard the Euclidean point approximation and use the true non-Euclidean arc costs of the $C$ matrix for the remaining steps of the ACHCI algorithm.
  3. 3.
    Find consecutive nodes $v_{i},v_{j}\in V$ in the subtour and $v_{k}\in V$ not in the subtour, that minimizes non-Euclidean insertion cost ratio $(C_{ik}+C_{kj})/C_{ij}$.
  4. 4.
    Insert $v_{k}$ between $v_{i}$ and $v_{j}$, updating the subtour.
  5. 5.
    Repeat Step 2 and Step 3, to obtain a Hamiltonian cycle.
For the example set of 25 non-Euclidean points, the subtour is initiated using the convex hull boundary points of $\widetilde{X}$ as shown in Fig. LABEL:Im1c:_CHullInit, after which the Euclidean coordinate approximation is discarded. The insertion cost ratios of Step 3, 4 and 5 use the true non-Euclidean cost matrix $C$ to obtain the complete tour, which is shown in Fig. LABEL:Im1d:_ACHCIsol.
4 Benchmark algorithms
Two heuristic algorithms and two metaheuristic algorithms are chosen as benchmarks for comparison with the ACHCI algorithm. The NN and NI heuristic algorithms are greedy algorithms based on selection rules and are known to perform well in constrained TSP applications[22, 23, 24]. Among metaheuristic algorithms that are commonly seen in TSP literature, the Genetic Algorithm (GA) and Ant Colony Optimization (ACO) are chosen as benchmarks because of their ability to utilize and update a population of solutions, offering a broader exploration of the solution space [25].
4.1 Nearest Neighbor Heuristic
Starting from a predefined or randomly selected initial node, the NN algorithm assigns the unvisited node associated with the lowest cost as the next node. This is repeated until all nodes are contained in the tour after which the starting location is visited again. The NN heuristic is as described below:
  1. 1.
    Initiate the subtour as a starting location
  2. 2.
    Append the subtour with the location with lowest non-Euclidean cost with respect to the location that is at the end of the current subtour
  3. 3.
    Repeat step 2 until all locations have been included
  4. 4.
    Return to the starting location
4.2 Nearest Insertion Heuristic
In contrast with the NN heuristic that is initiated with a single location, the NI heuristic starts with a subtour of 2 nodes and expands it by inserting unvisited nodes in increasing order of insertion cost:
  1. 1.
    Initialize a subtour as a starting location to itself.
  2. 2.
    Find consecutive nodes $(v_{i},v_{j})$ in the subtour, and $v_{k}$ not in the subtour, that minimizes non-Euclidean insertion cost $C_{ik}+C_{kj}-C_{ij}$.
  3. 3.
    Insert $v_{k}$ between $v_{i}$ and $v_{j}$, updating the subtour.
  4. 4.
    Repeat Step 2 and Step 3, to obtain a Hamiltonian cycle.
4.3 Genetic Algorithm
The GA is a popular metaheuristic algorithm, inspired by the mechanisms of natural selection and genetics [26]. The GA process is initialized with a population of $m$ randomly generated tours, known as chromosomes. The quality of each chromosome is defined by its fitness function, assigned as the inverse of the TSP tour length. In subsequent iterations, these chromosomes are modified to find solutions with higher fitness value. At any iteration, the incumbent solution is defined as the chromosome with the highest fitness values or lowest tour cost found up to that iteration. To ensure that the incumbent solution does not degrade as the iterations proceed, a defined number of high fitness chromosomes, called the elite population are carried over to the next generation without modification.
In each iteration, new tours called offspring are created by modifying some selected parent chromosomes of the current population. To ensure that promising parents have a higher chance of being selected and contributing to the next population, each chromosome $j$ with a fitness value $f_{j}$ is assigned a fitness proportionate probability given by $P_{j}=f_{j}/\sum_{i=1}^{m}f_{i}$. These probabilities are used by a Roulette Wheel Selection method to sample high fitness parents for the next generation [27].
The elite population and the Roulette Wheel selected chromosomes form the $m$ parents for the next generation. Of these parents, a fraction $p_{1}$ is randomly selected for the Ordered Crossover operation, where segments of two parent chromosomes are exchanged to produce two offspring. To ensure that duplicate cities do not feature in the resulting offspring, each offspring is constructed by choosing a sub-sequence of one parent while preserving the relative order of remaining cities of the other [28]. Next, a fraction $p_{2}$ of the resulting population is randomly sampled for the process of Swap Mutation whereby two random locations within a chromosome are selected and swapped to form a new offspring [29]. Of the resulting population, a fraction $p_{3}$ is then sampled for the Inversion Mutation, where a subtour segment of a parent chromosome is randomly chosen and its order is reversed [27].
For a predefined time limit, the operations of selection, crossover, and mutation iteratively explore the search space of TSP solutions to improve the incumbent solution. Upon termination, the incumbent solution is provided as the output tour.
4.4 Ant Colony Optimization
The ACO is another nature-inspired metaheuristic algorithm that mimics how a colony of ants finds the shortest paths to food sources based on distance and the strength of pheromone trails laid by individual ants [30]. While the GA uses stochastic processes to vary tour costs without explicitly evaluating edges in the arc set $A$, the ACO assigns pheromone strengths and heuristic values to the edges to inform exploration.
The ACO is initialized by computing local heuristic values $\eta_{ij}$ for each arc $(v_{i},v_{j})\in A$, assigned as the inverse of the cost $c(v_{i},v_{j})$ for $v_{i}\neq v_{j}$, and zero otherwise. The pheromone value $\tau_{ij}~{}\forall~{}(v_{i},v_{j})\in A$ is initially assigned a negligible value called the primary tracing, and in subsequent iterations, $\tau_{ij}$ will be updated to reflect the effect of edge $(v_{i},v_{j})$ on tour costs.
In each iteration, ants start from a randomly chosen location and build tours by sampling unvisited locations based on the transition probability $P_{ij}$ of arcs $(v_{i},v_{j})\in A$ [31]:
$P_{ij}=\frac{\tau_{ij}^{\alpha}\eta_{ij}^{\beta}}{\sum_{k\notin subtour}\tau_{ik}^{\alpha}\eta_{ik}^{\beta}}$ (4)
The influence of pheromone and heuristic information on tour edge selection is controlled by exponential parameters $\alpha$ and $\beta$ respectively. Let $c_{a}$ represent the cost of the tour sampled by ant $a$ of colony $\mathcal{A}$, obtained by sampling locations based on Eq.(4). After tours have been constructed for the ant colony, pheromone levels $\tau_{ij}$ on each edge $(v_{i},v_{j})$ are updated before the next iteration:
$\tau_{ij}\leftarrow(1-\rho)\tau_{ij}+\sum_{a\in\mathcal{A}}c_{a}^{-1}$ (5)
The evaporation rate parameter $\rho\in(0,1)$ discounts $\tau_{ij}$ in the next iteration to encourage exploration.
The iterative update of transition probabilities and pheromone trails results in an increased exploration of tours that exploit edges associated with lower tour costs. This guides the algorithm in finding improved tours based on the likelihood of an edge constituting the optimal tour, while the evaporation of pheromones encourages exploration of the entire search space to prevent premature convergence. The algorithm is terminated based on a convergence criteria such as the computation time limit, and the incumbent tour is provided as the output.
5 Computational Experiments
Figure 3: Two types of non-Euclidean modifications on the TSPLIB instance ‘t70’, shown with their ACHCI tours.
5.1 Non-Euclidean Test Cases
To compare the effectiveness of the ACHCI algorithms with the 4 benchmark algorithms, sufficiently diverse test instances were not readily available for the non-Euclidean TSP. For this reason, the popular TSPLIB instances [21] are modified to create non-Euclidean point clouds. In the first type of non-Euclidean modification, the $\mathcal{L}_{1}$ or Manhattan norm is defined as the cost between two locations in a TSPLIB point cloud [5]. In the second case, impassable separators are added as environmental constraints which act as obstacles.
The following steps ensure the reproducibility of adding separators to the TSPLIB point cloud:
  1. 1.
    Load a TSPLIB point cloud with 2D Cartesian coordinates.
  2. 2.
    Find the centroid of the point cloud.
  3. 3.
    Sort the points in order of increasing distance from the centroid.
  4. 4.
    Draw a line segment from the centroid to the farthest point. Trim its length by 5% from both ends. This line segment functions as the impassable separator.
  5. 5.
    For $k$ equiangular separators, rotate copies of the separator obtained in Step 4 about the centroid by multiples of $2\pi/k$ radians.
The ACHCI solution to the ‘t70’ instance outfitted with the $\mathcal{L}_{1}$ norm is shown in Fig. LABEL:t70_L1, where it can be seen that paths are only permissible along the $x$ and $y$ directions. A deviation from the Euclidean straight line path is also seen in the case with 4 added separators, as shown in Fig. LABEL:Im_4a:_NETSP. The distortion of the Euclidean space caused by non-Euclidean cost functions is quantified in literature [3, 32] by the deviation factor (DF):
$DF={|V|\choose 2}^{-1}\sum_{\begin{subarray}{c}(v_{i},v_{j})\in A\\ v_{i}\neq v_{j}\end{subarray}}\frac{\Delta(v_{i},v_{j})}{\delta(v_{i},v_{j})}$ (6)
In Eq. (6), for each pair of points $(v_{i},v_{j})\in A$, the Euclidean distance is denoted by $\delta(v_{i},v_{j})$ while $\Delta(v_{i},v_{j})$ is the true non-Euclidean cost. If this cost is defined by the $\mathcal{L}_{1}$ norm, then $\Delta(v_{i},v_{j})=|x_{i_{1}}-x_{j_{1}}|+|x_{i_{2}}-x_{j_{2}}|$, where $(x_{i_{1}},x_{i_{2}})$ and $(x_{j_{1}},x_{j_{2}})$ define the coordinates of $v_{i}$ and $v_{j}$ respectively. When instead, obstacles like impassable separators cause the deviation from the straight light path, the shortest true path length between pairs of points in $(v_{i},v_{j})\in A$ is computed using Dijkstra’s shortest path algorithm [33] to find $\Delta(v_{i},v_{j})$.
Table 1: Metaheuristic hyperparameter settings
Genetic Algorithm
Population Size: 1000
Elite Count: 10
Ordered Crossover Fraction $p_{1}$: 0.9
Swap Mutation Fraction $p_{2}$: 0.02
Inversion Fraction $p_{3}$: 0.07
Ant Colony Optimization
Colony Size 1000
Primary tracing: $10^{-4}$
Exponential Pheromone Parameter $\beta$: 3
Exponential Heuristic Parameter $\alpha$: 5
Evaporation Coefficient $\rho$: 0.15
5.2 Performance Analysis
Figure 4: Convergence of metaheuristic algorithms
To analyze the performance of the ACHCI heuristic, extensive computational experiments were conducted in a Matlab R2022a environment on an Intel Xeon E5-2680 v4 CPU clocked at 2.4 GHz at the Ohio Super Computer [34]. The utility of the approximate algorithms compared in this paper lies in their ability to obtain solutions quickly, and for this reason, each algorithm is executed up to a time limit of 60 seconds. The hyperparameters that define the metaheuristic algorithms influence the convergence of incumbent solutions. Their values are listed in Table 1. To account for the stochasticity involved in the GA and ACO algorithms, each algorithm is repeated 28 times for each TSPLIB case and the distribution of solutions is noted upon termination at 60 seconds.
The convergence of GA and ACO incumbent solutions along with their 95% confidence interval is shown in Fig. LABEL:Im:roC100convergence for the case where the $\mathcal{L}_{1}$ norm is applied to the TSPLIB instance ‘roC100’ with 100 points. The GA only uses tour cost as a metric for tour improvement which results in rapid rates of convergence. Contrarily, because the ACO samples locations and updates pheromone values $\tau_{ij}$ of each edge as per Eq. (4) and (5) respectively, the computational demand per iteration is higher. While this results in a slower convergence rate, its incumbent solution costs are initially lower than that of GA. In Fig. LABEL:Im:roB150convergence where 4 separators were added to the 150 points of instance ‘roB150’, it is seen that despite the larger number of points, the initial performance of the ACO is still superior.
Performance comparisons are conducted for 57 TSPLIB instances, with the number of points varying from 51 to 1,400. For each TSPLIB instance, one non-Euclidean case is created by assigning costs using the $\mathcal{L}_{1}$ norm, and 3 cases are created by adding 4, 16 or 64 impassable separators. For the sake of brevity, only 14 instances are tabulated in Table 2, though an unabridged version can be found in [35].
Table 2: TSP-NE Solution Costs
TSPLIB DF   (Cause) ACHCI NN NI GA $\mu(\sigma)$ ACO $\mu(\sigma)$
eil51 1.28 $(\mathcal{L}_{1}norm)$ 5.76e+02 6.62e+02 6.16e+02 5.75e+02 (1.82e+00) 5.69e+02 (8.36e+00)
1.18 $(4~{}separators)$ 4.83e+02 5.60e+02 5.54e+02 4.90e+02 (1.44e+01) 4.86e+02 (1.05e+01)
1.42 $(16~{}separators)$ 7.33e+02 7.91e+02 8.20e+02 7.30e+02 (9.95e+00) 7.40e+02 (1.00e+01)
1.55 $(64~{}separators)$ 1.31e+03 1.35e+03 1.33e+03 1.31e+03 (2.20e+00) 1.37e+03 (1.34e+01)
erlin52 1.27 $(\mathcal{L}_{1}norm)$ 9.92e+03 1.12e+04 1.08e+04 9.93e+03 (2.32e+02) 9.89e+03 (4.01e+01)
1.14 $(4~{}separators)$ 8.98e+03 9.96e+03 9.81e+03 8.97e+03 (6.06e+00) 9.02e+03 (5.98e+01)
1.30 $(16~{}separators)$ 1.37e+04 1.55e+04 1.50e+04 1.36e+04 (3.30e+01) 1.41e+04 (1.94e+02)
1.43 $(64~{}separators)$ 2.16e+04 2.33e+04 2.25e+04 2.16e+04 (2.73e+01) 2.28e+04 (5.32e+01)
t70 1.27 $(\mathcal{L}_{1}norm)$ 8.60e+02 1.00e+03 9.32e+02 8.92e+02 (1.77e+01) 9.26e+02 (3.58e+01)
1.21 $(4~{}separators)$ 7.69e+02 8.86e+02 8.88e+02 8.17e+02 (6.81e+01) 7.86e+02 (3.68e+01)
1.44 $(16~{}separators)$ 1.19e+03 1.23e+03 1.28e+03 1.19e+03 (1.85e+00) 1.18e+03 (5.96e+00)
1.60 $(64~{}separators)$ 2.06e+03 2.22e+03 2.18e+03 2.06e+03 (4.62e+00) 2.16e+03 (3.96e+01)
eil76 1.28 $(\mathcal{L}_{1}norm)$ 7.50e+02 8.74e+02 7.82e+02 7.50e+02 (3.51e+00) 7.57e+02 (8.30e+00)
1.19 $(4~{}separators)$ 6.27e+02 7.05e+02 6.99e+02 6.23e+02 (5.26e+00) 6.34e+02 (3.23e+01)
1.40 $(16~{}separators)$ 9.17e+02 9.48e+02 9.75e+02 9.12e+02 (4.06e+00) 8.88e+02 (1.24e+01)
1.52 $(64~{}separators)$ 1.61e+03 1.69e+03 1.69e+03 1.61e+03 (1.47e+00) 1.72e+03 (4.17e+01)
pr76 1.26 $(\mathcal{L}_{1}norm)$ 1.46e+05 1.78e+05 1.55e+05 1.45e+05 (6.62e+02) 1.44e+05 (8.55e+02)
1.19 $(4~{}separators)$ 1.22e+05 1.53e+05 1.38e+05 1.30e+05 (1.32e+04) 1.22e+05 (1.33e+03)
1.46 $(16~{}separators)$ 1.85e+05 1.99e+05 2.19e+05 1.85e+05 (2.20e+02) 1.92e+05 (2.33e+03)
1.62 $(64~{}separators)$ 3.72e+05 3.97e+05 3.76e+05 3.72e+05 (3.47e+02) 4.00e+05 (3.13e+03)
rat99 1.24 $(\mathcal{L}_{1}norm)$ 1.63e+03 1.84e+03 1.86e+03 1.65e+03 (6.65e+01) 1.70e+03 (5.32e+01)
1.15 $(4~{}separators)$ 1.36e+03 1.57e+03 1.55e+03 1.40e+03 (1.16e+02) 1.49e+03 (8.25e+01)
1.38 $(16~{}separators)$ 2.08e+03 2.35e+03 2.35e+03 2.09e+03 (7.87e+01) 2.09e+03 (6.16e+01)
1.54 $(64~{}separators)$ 4.03e+03 4.20e+03 4.28e+03 4.03e+03 (3.50e+00) 4.23e+03 (4.25e+01)
roC100 1.25 $(\mathcal{L}_{1}norm)$ 2.73e+04 3.13e+04 3.17e+04 2.97e+04 (2.39e+03) 3.04e+04 (6.06e+02)
1.17 $(4~{}separators)$ 2.36e+04 2.87e+04 2.82e+04 2.56e+04 (3.33e+03) 2.63e+04 (1.68e+03)
1.41 $(16~{}separators)$ 3.74e+04 4.04e+04 4.19e+04 3.74e+04 (6.68e+01) 3.81e+04 (5.33e+02)
1.61 $(64~{}separators)$ 7.14e+04 7.83e+04 7.48e+04 7.21e+04 (1.63e+03) 7.72e+04 (5.90e+01)
roD100 1.25 $(\mathcal{L}_{1}norm)$ 2.69e+04 3.26e+04 3.16e+04 2.85e+04 (7.79e+02) 3.02e+04 (8.12e+02)
1.18 $(4~{}separators)$ 2.36e+04 2.60e+04 2.77e+04 2.47e+04 (1.03e+03) 2.58e+04 (2.34e+02)
1.37 $(16~{}separators)$ 3.39e+04 3.87e+04 4.09e+04 3.37e+04 (3.37e+02) 3.59e+04 (1.10e+03)
1.56 $(64~{}separators)$ 6.66e+04 7.17e+04 7.33e+04 6.66e+04 (1.57e+03) 7.08e+04 (1.36e+03)
roE100 1.25 $(\mathcal{L}_{1}norm)$ 2.87e+04 3.44e+04 3.20e+04 3.01e+04 (1.92e+03) 3.09e+04 (3.76e+02)
1.19 $(4~{}separators)$ 2.36e+04 2.79e+04 2.88e+04 2.59e+04 (4.81e+03) 2.61e+04 (7.11e+02)
1.39 $(16~{}separators)$ 3.72e+04 3.88e+04 4.20e+04 3.71e+04 (8.00e+01) 3.67e+04 (1.49e+02)
1.61 $(64~{}separators)$ 6.53e+04 7.13e+04 7.11e+04 6.59e+04 (1.34e+03) 7.07e+04 (1.28e+03)
roA100 1.24 $(\mathcal{L}_{1}norm)$ 2.73e+04 3.17e+04 3.26e+04 2.86e+04 (8.06e+02) 3.00e+04 (1.50e+03)
1.15 $(4~{}separators)$ 2.37e+04 2.84e+04 2.79e+04 2.41e+04 (1.42e+03) 2.56e+04 (6.70e+01)
1.37 $(16~{}separators)$ 3.65e+04 3.94e+04 4.13e+04 3.65e+04 (3.14e+02) 3.70e+04 (5.09e+02)
1.56 $(64~{}separators)$ 7.22e+04 7.45e+04 7.52e+04 7.21e+04 (8.18e+01) 7.64e+04 (4.39e+02)
roB100 1.25 $(\mathcal{L}_{1}norm)$ 2.95e+04 3.57e+04 3.23e+04 3.08e+04 (2.27e+03) 3.14e+04 (8.75e+02)
1.18 $(4~{}separators)$ 2.68e+04 2.98e+04 2.85e+04 2.68e+04 (2.65e+02) 2.90e+04 (3.14e+02)
1.41 $(16~{}separators)$ 3.81e+04 4.02e+04 4.31e+04 3.84e+04 (1.22e+03) 3.80e+04 (2.09e+02)
1.62 $(64~{}separators)$ 7.68e+04 8.02e+04 8.20e+04 7.66e+04 (1.64e+02) 8.10e+04 (1.72e+03)
rd100 1.27 $(\mathcal{L}_{1}norm)$ 1.07e+04 1.23e+04 1.18e+04 1.07e+04 (1.10e+01) 1.11e+04 (4.72e+02)
1.21 $(4~{}separators)$ 9.02e+03 9.98e+03 1.06e+04 9.34e+03 (3.43e+02) 9.79e+03 (4.19e+01)
1.48 $(16~{}separators)$ 1.36e+04 1.34e+04 1.61e+04 1.35e+04 (2.09e+01) 1.37e+04 (2.93e+02)
1.70 $(64~{}separators)$ 2.86e+04 2.99e+04 2.95e+04 2.86e+04 (2.59e+01) 3.06e+04 (5.62e+02)
eil101 1.28 $(\mathcal{L}_{1}norm)$ 8.82e+02 9.76e+02 9.10e+02 8.82e+02 (4.21e+00) 9.50e+02 (1.53e+01)
1.21 $(4~{}separators)$ 7.22e+02 8.65e+02 7.59e+02 7.34e+02 (3.06e+01) 7.83e+02 (3.06e+01)
1.48 $(16~{}separators)$ 1.02e+03 1.11e+03 1.16e+03 1.03e+03 (2.30e+01) 1.03e+03 (4.35e+01)
1.68 $(64~{}separators)$ 2.07e+03 2.13e+03 2.18e+03 2.06e+03 (1.67e+00) 2.20e+03 (1.78e+01)
in105 1.23 $(\mathcal{L}_{1}norm)$ 1.71e+04 1.94e+04 2.10e+04 1.84e+04 (1.61e+03) 1.80e+04 (5.70e+02)
1.08 $(4~{}separators)$ 1.68e+04 1.95e+04 1.86e+04 1.69e+04 (9.56e+02) 1.82e+04 (3.75e+02)
1.38 $(16~{}separators)$ 2.34e+04 2.55e+04 2.79e+04 2.34e+04 (5.82e+01) 2.46e+04 (2.08e+00)
1.63 $(64~{}separators)$ 4.25e+04 4.41e+04 4.73e+04 4.22e+04 (2.26e+02) 4.45e+04 (7.86e+02)
Table 2 (contd.): TSP-NE Solution Costs
TSPLIB DF   (Cause) ACHCI NN NI GA $\mu(\sigma)$ ACO $\mu(\sigma)$
pr107 1.17 $(\mathcal{L}_{1}norm)$ 5.18e+04 5.79e+04 5.78e+04 5.28e+04 (1.77e+03) 5.16e+04 (2.58e+02)
1.03 $(4~{}separators)$ 4.64e+04 5.39e+04 5.19e+04 4.91e+04 (4.20e+03) 4.93e+04 (1.26e+03)
1.23 $(16~{}separators)$ 5.41e+04 6.26e+04 6.23e+04 5.58e+04 (2.55e+03) 5.83e+04 (4.51e+02)
1.47 $(64~{}separators)$ 8.22e+04 8.86e+04 8.66e+04 8.43e+04 (4.62e+02) 8.74e+04 (6.62e+02)
pr124 1.26 $(\mathcal{L}_{1}norm)$ 7.14e+04 8.60e+04 8.36e+04 8.05e+04 (6.22e+03) 7.26e+04 (1.91e+03)
1.20 $(4~{}separators)$ 6.85e+04 8.66e+04 8.38e+04 7.13e+04 (6.40e+03) 7.37e+04 (2.45e+03)
1.48 $(16~{}separators)$ 1.10e+05 1.12e+05 1.20e+05 1.10e+05 (1.87e+02) 1.10e+05 (2.33e+03)
1.77 $(64~{}separators)$ 2.30e+05 2.41e+05 2.41e+05 2.34e+05 (6.89e+03) 2.52e+05 (1.48e+03)
bier127 1.28 $(\mathcal{L}_{1}norm)$ 1.60e+05 1.80e+05 1.81e+05 1.68e+05 (7.68e+03) 1.67e+05 (2.63e+03)
1.16 $(4~{}separators)$ 1.42e+05 1.76e+05 1.54e+05 1.42e+05 (7.20e+03) 1.59e+05 (5.69e+03)
1.31 $(16~{}separators)$ 1.86e+05 2.06e+05 2.11e+05 1.92e+05 (1.68e+04) 2.11e+05 (5.72e+03)
1.45 $(64~{}separators)$ 3.80e+05 4.27e+05 3.93e+05 3.85e+05 (6.03e+03) 4.35e+05 (5.68e+03)
h130 1.28 $(\mathcal{L}_{1}norm)$ 8.14e+03 1.01e+04 9.28e+03 8.50e+03 (4.27e+02) 8.79e+03 (1.36e+02)
1.19 $(4~{}separators)$ 6.90e+03 8.01e+03 7.64e+03 7.13e+03 (7.22e+02) 7.37e+03 (1.58e+02)
1.45 $(16~{}separators)$ 9.99e+03 9.98e+03 1.13e+04 1.00e+04 (3.65e+02) 9.99e+03 (8.62e+01)
1.65 $(64~{}separators)$ 2.13e+04 2.18e+04 2.26e+04 2.21e+04 (1.20e+03) 2.26e+04 (2.55e+02)
pr136 1.25 $(\mathcal{L}_{1}norm)$ 1.29e+05 1.20e+05 1.34e+05 1.31e+05 (4.01e+03) 1.33e+05 (2.35e+03)
1.19 $(4~{}separators)$ 1.17e+05 1.38e+05 1.24e+05 1.21e+05 (9.21e+03) 1.30e+05 (6.21e+03)
1.44 $(16~{}separators)$ 1.61e+05 1.63e+05 1.70e+05 1.61e+05 (2.02e+03) 1.64e+05 (2.56e+03)
1.66 $(64~{}separators)$ 3.02e+05 2.98e+05 3.24e+05 3.15e+05 (2.24e+04) 3.21e+05 (4.35e+03)
pr144 1.24 $(\mathcal{L}_{1}norm)$ 7.36e+04 7.66e+04 8.13e+04 7.87e+04 (8.92e+03) 7.27e+04 (2.88e+02)
1.17 $(4~{}separators)$ 6.82e+04 6.83e+04 7.98e+04 7.39e+04 (9.32e+03) 6.99e+04 (7.85e+02)
1.33 $(16~{}separators)$ 9.43e+04 1.02e+05 1.10e+05 9.88e+04 (8.50e+03) 1.00e+05 (4.33e+03)
1.57 $(64~{}separators)$ 1.78e+05 1.87e+05 1.90e+05 1.87e+05 (8.40e+03) 1.91e+05 (2.86e+03)
roB150 1.25 $(\mathcal{L}_{1}norm)$ 3.53e+04 4.19e+04 3.83e+04 3.97e+04 (6.10e+03) 3.96e+04 (3.49e+02)
1.17 $(4~{}separators)$ 3.03e+04 3.52e+04 3.38e+04 3.39e+04 (4.07e+03) 3.38e+04 (1.01e+03)
1.40 $(16~{}separators)$ 4.35e+04 4.61e+04 4.76e+04 4.66e+04 (3.44e+03) 4.49e+04 (8.14e+02)
1.62 $(64~{}separators)$ 8.31e+04 8.81e+04 8.94e+04 9.73e+04 (1.66e+04) 8.77e+04 (1.55e+03)
roA150 1.25 $(\mathcal{L}_{1}norm)$ 3.58e+04 4.23e+04 3.94e+04 4.03e+04 (2.29e+03) 3.95e+04 (1.17e+03)
1.17 $(4~{}separators)$ 2.93e+04 3.69e+04 3.27e+04 3.28e+04 (2.69e+03) 3.35e+04 (1.93e+03)
1.39 $(16~{}separators)$ 4.16e+04 4.44e+04 4.55e+04 4.50e+04 (7.58e+03) 4.39e+04 (1.07e+03)
1.59 $(64~{}separators)$ 8.92e+04 9.50e+04 9.36e+04 9.70e+04 (1.03e+04) 9.53e+04 (1.62e+03)
h150 1.27 $(\mathcal{L}_{1}norm)$ 8.71e+03 9.93e+03 1.01e+04 1.02e+04 (9.87e+02) 9.74e+03 (3.49e+02)
1.16 $(4~{}separators)$ 7.05e+03 8.79e+03 8.37e+03 8.31e+03 (8.87e+02) 7.80e+03 (7.46e+02)
1.35 $(16~{}separators)$ 1.06e+04 1.00e+04 1.14e+04 1.19e+04 (1.98e+03) 1.05e+04 (2.90e+01)
1.50 $(64~{}separators)$ 2.00e+04 2.13e+04 2.13e+04 2.24e+04 (1.94e+03) 2.17e+04 (4.61e+02)
pr152 1.22 $(\mathcal{L}_{1}norm)$ 9.20e+04 9.64e+04 1.05e+05 9.58e+04 (1.34e+04) 9.70e+04 (3.23e+03)
1.09 $(4~{}separators)$ 7.93e+04 9.85e+04 9.22e+04 8.54e+04 (1.96e+03) 8.66e+04 (4.66e+03)
1.31 $(16~{}separators)$ 1.10e+05 1.32e+05 1.21e+05 1.13e+05 (6.32e+03) 1.18e+05 (2.52e+03)
1.46 $(64~{}separators)$ 1.82e+05 1.90e+05 1.92e+05 1.92e+05 (1.06e+04) 1.92e+05 (9.75e+02)
u159 1.24 $(\mathcal{L}_{1}norm)$ 5.58e+04 5.96e+04 6.16e+04 6.45e+04 (1.25e+04) 6.00e+04 (1.81e+03)
1.13 $(4~{}separators)$ 5.09e+04 5.49e+04 5.67e+04 5.75e+04 (7.10e+03) 5.60e+04 (1.54e+03)
1.40 $(16~{}separators)$ 7.22e+04 7.33e+04 7.79e+04 8.02e+04 (8.72e+03) 7.35e+04 (4.83e+03)
1.63 $(64~{}separators)$ 1.38e+05 1.37e+05 1.43e+05 1.56e+05 (1.95e+04) 1.42e+05 (7.34e+02)
rat195 1.24 $(\mathcal{L}_{1}norm)$ 3.31e+03 3.62e+03 3.52e+03 5.54e+03 (9.97e+02) 3.47e+03 (1.00e+02)
1.15 $(4~{}separators)$ 2.67e+03 3.10e+03 2.92e+03 4.57e+03 (6.53e+02) 3.04e+03 (7.75e+00)
1.38 $(16~{}separators)$ 3.45e+03 3.68e+03 3.74e+03 7.29e+03 (1.09e+03) 3.73e+03 (8.31e+01)
1.54 $(64~{}separators)$ 7.77e+03 7.67e+03 8.31e+03 1.16e+04 (9.04e+02) 7.93e+03 (8.93e+01)
d198 1.27 $(\mathcal{L}_{1}norm)$ 2.00e+04 2.24e+04 2.14e+04 2.86e+04 (7.41e+03) 2.18e+04 (7.32e+02)
1.11 $(4~{}separators)$ 1.77e+04 1.84e+04 1.85e+04 2.41e+04 (4.36e+03) 1.90e+04 (9.05e+01)
1.51 $(16~{}separators)$ 2.05e+04 2.33e+04 2.21e+04 3.02e+04 (5.36e+03) 2.29e+04 (6.54e+02)
1.99 $(64~{}separators)$ 4.13e+04 4.28e+04 4.26e+04 5.82e+04 (1.53e+04) 4.30e+04 (4.43e+02)
roA200 1.25 $(\mathcal{L}_{1}norm)$ 4.06e+04 4.53e+04 4.48e+04 7.97e+04 (2.14e+04) 4.44e+04 (1.47e+03)
1.18 $(4~{}separators)$ 3.45e+04 4.18e+04 3.71e+04 6.97e+04 (1.40e+04) 3.93e+04 (6.54e+02)
1.42 $(16~{}separators)$ 4.60e+04 5.28e+04 5.12e+04 1.05e+05 (1.58e+04) 4.96e+04 (7.63e+02)
1.62 $(64~{}separators)$ 9.66e+04 9.82e+04 1.07e+05 1.68e+05 (1.59e+04) 1.02e+05 (9.58e+02)
Table 2 (contd.): TSP-NE Solution Costs
TSPLIB DF   (Cause) ACHCI NN NI GA $\mu(\sigma)$ ACO $\mu(\sigma)$
roB200 1.25 $(\mathcal{L}_{1}norm)$ 4.01e+04 4.77e+04 4.55e+04 8.87e+04 (1.35e+04) 4.52e+04 (3.75e+02)
1.17 $(4~{}separators)$ 3.41e+04 3.80e+04 3.80e+04 6.56e+04 (1.19e+04) 3.80e+04 (2.01e+03)
1.40 $(16~{}separators)$ 4.67e+04 5.47e+04 5.12e+04 9.97e+04 (1.99e+04) 5.13e+04 (1.31e+03)
1.60 $(64~{}separators)$ 1.07e+05 1.08e+05 1.13e+05 1.71e+05 (1.55e+04) 1.11e+05 (1.06e+03)
ts225 1.26 $(\mathcal{L}_{1}norm)$ 1.61e+05 1.60e+05 1.81e+05 5.37e+05 (7.56e+04) 1.57e+05 (3.84e+03)
1.17 $(4~{}separators)$ 1.61e+05 1.57e+05 1.94e+05 5.16e+05 (1.54e+05) 1.61e+05 (1.67e+03)
1.42 $(16~{}separators)$ 3.80e+05 3.84e+05 3.86e+05 8.60e+05 (7.77e+04) 3.96e+05 (1.07e+04)
1.54 $(64~{}separators)$ 6.84e+05 6.69e+05 6.95e+05 1.10e+06 (7.44e+04) 7.72e+05 (1.98e+04)
tsp225 1.26 $(\mathcal{L}_{1}norm)$ 4.55e+03 5.66e+03 4.99e+03 1.31e+04 (2.86e+03) 4.96e+03 (8.74e+01)
1.19 $(4~{}separators)$ 4.48e+03 5.57e+03 4.75e+03 1.30e+04 (3.98e+03) 5.13e+03 (2.15e+02)
1.45 $(16~{}separators)$ 5.80e+03 6.74e+03 6.29e+03 1.81e+04 (2.62e+03) 6.49e+03 (2.41e+02)
1.64 $(64~{}separators)$ 1.40e+04 1.43e+04 1.50e+04 2.85e+04 (4.93e+03) 1.48e+04 (3.46e+02)
pr226 1.23 $(\mathcal{L}_{1}norm)$ 9.61e+04 1.13e+05 1.16e+05 3.69e+05 (1.28e+05) 1.06e+05 (9.36e+03)
1.19 $(4~{}separators)$ 9.13e+04 1.20e+05 1.10e+05 3.84e+05 (9.76e+04) 1.03e+05 (4.39e+03)
1.48 $(16~{}separators)$ 1.44e+05 1.53e+05 1.66e+05 5.73e+05 (1.59e+05) 1.61e+05 (3.55e+03)
1.87 $(64~{}separators)$ 2.85e+05 2.98e+05 3.03e+05 8.84e+05 (2.32e+05) 3.12e+05 (2.45e+03)
gil262 1.27 $(\mathcal{L}_{1}norm)$ 3.17e+03 3.67e+03 3.51e+03 1.31e+04 (3.14e+03) 3.65e+03 (1.77e+02)
1.22 $(4~{}separators)$ 2.87e+03 3.20e+03 3.05e+03 1.27e+04 (2.69e+03) 3.17e+03 (5.65e+01)
1.50 $(16~{}separators)$ 3.66e+03 3.97e+03 4.05e+03 1.81e+04 (2.34e+03) 4.01e+03 (1.34e+02)
1.70 $(64~{}separators)$ 8.58e+03 8.53e+03 9.28e+03 2.31e+04 (1.70e+03) 9.02e+03 (9.49e+01)
pr264 1.19 $(\mathcal{L}_{1}norm)$ 6.12e+04 6.06e+04 6.64e+04 2.64e+05 (7.30e+04) 6.21e+04 (3.30e+03)
1.03 $(4~{}separators)$ 5.48e+04 6.03e+04 6.00e+04 2.16e+05 (5.41e+04) 6.14e+04 (1.80e+03)
1.50 $(16~{}separators)$ 6.51e+04 7.01e+04 7.54e+04 4.05e+05 (1.04e+05) 7.40e+04 (2.16e+03)
2.00 $(64~{}separators)$ 1.18e+05 1.28e+05 1.25e+05 6.65e+05 (1.15e+05) 1.27e+05 (3.80e+03)
pr299 1.23 $(\mathcal{L}_{1}norm)$ 6.40e+04 6.92e+04 6.93e+04 3.86e+05 (4.25e+04) 7.10e+04 (2.54e+03)
1.22 $(4~{}separators)$ 5.48e+04 6.00e+04 6.01e+04 3.85e+05 (8.54e+04) 6.45e+04 (1.61e+03)
1.49 $(16~{}separators)$ 6.82e+04 7.80e+04 7.69e+04 4.56e+05 (9.63e+04) 8.23e+04 (1.16e+03)
1.79 $(64~{}separators)$ 1.64e+05 1.63e+05 1.75e+05 6.20e+05 (4.02e+04) 1.75e+05 (5.08e+03)
in318 1.26 $(\mathcal{L}_{1}norm)$ 5.40e+04 6.00e+04 5.99e+04 3.89e+05 (7.69e+04) 6.07e+04 (1.57e+03)
1.21 $(4~{}separators)$ 4.92e+04 5.57e+04 5.30e+04 3.47e+05 (3.91e+04) 5.64e+04 (7.37e+02)
1.46 $(16~{}separators)$ 6.42e+04 6.96e+04 7.02e+04 4.37e+05 (5.57e+04) 7.11e+04 (2.23e+03)
1.69 $(64~{}separators)$ 1.40e+05 1.44e+05 1.53e+05 5.43e+05 (3.81e+04) 1.52e+05 (2.09e+03)
in318 1.26 $(\mathcal{L}_{1}norm)$ 5.40e+04 6.00e+04 5.99e+04 3.87e+05 (7.16e+04) 6.03e+04 (1.13e+03)
1.21 $(4~{}separators)$ 4.92e+04 5.57e+04 5.30e+04 3.48e+05 (4.80e+04) 5.66e+04 (7.92e+02)
1.46 $(16~{}separators)$ 6.42e+04 6.96e+04 7.02e+04 4.44e+05 (3.70e+04) 7.21e+04 (1.18e+03)
1.69 $(64~{}separators)$ 1.40e+05 1.44e+05 1.53e+05 5.28e+05 (9.64e+04) 1.51e+05 (2.32e+03)
rd400 1.27 $(\mathcal{L}_{1}norm)$ 2.09e+04 2.35e+04 2.37e+04 1.82e+05 (1.70e+04) 2.62e+04 (3.96e+03)
1.20 $(4~{}separators)$ 1.77e+04 2.10e+04 1.91e+04 1.65e+05 (1.48e+04) 2.20e+04 (2.44e+03)
1.45 $(16~{}separators)$ 2.18e+04 2.41e+04 2.26e+04 1.98e+05 (7.94e+03) 2.66e+04 (4.00e+03)
1.65 $(64~{}separators)$ 4.75e+04 4.35e+04 5.13e+04 2.18e+05 (1.04e+04) 5.24e+04 (1.02e+04)
fl417 1.19 $(\mathcal{L}_{1}norm)$ 1.47e+04 1.83e+04 1.79e+04 3.46e+05 (3.91e+04) 1.85e+04 (1.66e+03)
1.04 $(4~{}separators)$ 1.30e+04 1.63e+04 1.45e+04 3.13e+05 (3.04e+04) 1.62e+04 (1.44e+03)
1.22 $(16~{}separators)$ 2.00e+04 2.17e+04 2.21e+04 3.42e+05 (4.95e+04) 2.25e+04 (3.24e+03)
1.50 $(64~{}separators)$ 3.47e+04 3.81e+04 3.66e+04 3.79e+05 (3.75e+04) 4.06e+04 (3.42e+03)
pr439 1.28 $(\mathcal{L}_{1}norm)$ 1.44e+05 1.57e+05 1.56e+05 1.57e+06 (1.40e+05) 1.70e+05 (7.26e+03)
1.29 $(4~{}separators)$ 1.32e+05 1.50e+05 1.44e+05 1.54e+06 (1.43e+05) 1.68e+05 (3.63e+03)
1.60 $(16~{}separators)$ 1.70e+05 1.84e+05 1.83e+05 1.79e+06 (7.26e+04) 2.10e+05 (5.33e+03)
1.99 $(64~{}separators)$ 4.03e+05 4.02e+05 4.19e+05 2.09e+06 (9.10e+04) 4.56e+05 (3.84e+03)
pcb442 1.27 $(\mathcal{L}_{1}norm)$ 6.28e+04 7.58e+04 7.08e+04 6.87e+05 (4.61e+04) 8.13e+04 (1.34e+03)
1.22 $(4~{}separators)$ 5.71e+04 6.57e+04 6.34e+04 6.49e+05 (4.51e+04) 7.72e+04 (2.30e+03)
1.49 $(16~{}separators)$ 7.50e+04 8.11e+04 7.87e+04 7.55e+05 (3.18e+04) 9.56e+04 (2.94e+03)
1.71 $(64~{}separators)$ 1.77e+05 1.66e+05 1.91e+05 8.70e+05 (4.28e+04) 2.09e+05 (5.41e+03)
d493 1.28 $(\mathcal{L}_{1}norm)$ 4.79e+04 5.28e+04 5.20e+04 4.14e+05 (3.98e+04) 6.00e+04 (2.65e+03)
1.20 $(4~{}separators)$ 3.99e+04 4.87e+04 4.17e+04 3.71e+05 (1.28e+04) 5.18e+04 (1.70e+03)
1.48 $(16~{}separators)$ 4.73e+04 5.49e+04 4.84e+04 4.28e+05 (4.89e+04) 6.50e+04 (1.79e+03)
1.79 $(64~{}separators)$ 1.02e+05 1.16e+05 1.03e+05 4.86e+05 (6.96e+03) 1.32e+05 (4.20e+03)
Table 2 (contd.): TSP-NE Solution Costs
TSPLIB DF   (Cause) ACHCI NN NI GA $\mu(\sigma)$ ACO $\mu(\sigma)$
u574 1.26 $(\mathcal{L}_{1}norm)$ 5.07e+04 5.61e+04 5.40e+04 6.62e+05 (2.59e+04) -
1.19 $(4~{}separators)$ 4.25e+04 4.98e+04 4.67e+04 6.00e+05 (1.76e+04) -
1.46 $(16~{}separators)$ 5.07e+04 5.91e+04 5.40e+04 7.15e+05 (4.03e+04) -
1.68 $(64~{}separators)$ 1.18e+05 1.09e+05 1.22e+05 7.75e+05 (1.98e+04) -
rat575 1.25 $(\mathcal{L}_{1}norm)$ 9.55e+03 1.08e+04 1.02e+04 1.07e+05 (4.53e+03) -
1.17 $(4~{}separators)$ 7.64e+03 8.79e+03 8.12e+03 9.93e+04 (5.03e+03) -
1.40 $(16~{}separators)$ 8.81e+03 1.01e+04 9.10e+03 1.13e+05 (1.85e+03) -
1.60 $(64~{}separators)$ 1.95e+04 1.74e+04 1.95e+04 1.23e+05 (3.91e+03) -
p654 1.17 $(\mathcal{L}_{1}norm)$ 4.48e+04 5.41e+04 5.04e+04 1.86e+06 (1.50e+05) -
1.04 $(4~{}separators)$ 3.98e+04 5.09e+04 4.64e+04 1.66e+06 (1.33e+05) -
1.32 $(16~{}separators)$ 5.76e+04 5.96e+04 6.02e+04 1.86e+06 (7.27e+04) -
1.65 $(64~{}separators)$ 1.02e+05 1.06e+05 1.04e+05 2.02e+06 (1.25e+05) -
d657 1.27 $(\mathcal{L}_{1}norm)$ 6.90e+04 8.27e+04 7.45e+04 9.03e+05 (4.53e+04) -
1.23 $(4~{}separators)$ 5.60e+04 6.91e+04 6.05e+04 8.38e+05 (1.24e+04) -
1.54 $(16~{}separators)$ 6.63e+04 8.26e+04 6.75e+04 9.79e+05 (4.18e+04) -
1.79 $(64~{}separators)$ 1.53e+05 1.61e+05 1.57e+05 1.07e+06 (2.16e+04) -
u724 1.26 $(\mathcal{L}_{1}norm)$ 5.67e+04 6.50e+04 6.21e+04 9.03e+05 (2.29e+04) -
1.19 $(4~{}separators)$ 4.82e+04 5.62e+04 5.18e+04 8.42e+05 (1.75e+04) -
1.45 $(16~{}separators)$ 5.55e+04 6.47e+04 5.94e+04 9.60e+05 (1.65e+04) -
1.66 $(64~{}separators)$ 1.22e+05 1.13e+05 1.28e+05 1.03e+06 (2.42e+04) -
rat783 1.25 $(\mathcal{L}_{1}norm)$ 1.24e+04 1.40e+04 1.35e+04 1.91e+05 (3.47e+03) -
1.17 $(4~{}separators)$ 1.01e+04 1.16e+04 1.07e+04 1.72e+05 (9.45e+03) -
1.41 $(16~{}separators)$ 1.13e+04 1.34e+04 1.18e+04 1.96e+05 (6.14e+03) -
1.60 $(64~{}separators)$ 2.26e+04 2.15e+04 2.35e+04 2.09e+05 (2.26e+03) -
pr1002 1.27 $(\mathcal{L}_{1}norm)$ 3.54e+05 4.02e+05 3.78e+05 7.33e+06 (1.51e+05) -
1.23 $(4~{}separators)$ 2.93e+05 3.48e+05 3.13e+05 6.81e+06 (1.80e+05) -
1.53 $(16~{}separators)$ 3.53e+05 4.03e+05 3.59e+05 7.85e+06 (1.81e+05) -
1.80 $(64~{}separators)$ 7.33e+05 7.33e+05 7.66e+05 8.38e+06 (1.52e+05) -
u1060 1.24 $(\mathcal{L}_{1}norm)$ 3.04e+05 3.65e+05 - 7.45e+06 (1.62e+05) -
1.18 $(4~{}separators)$ 2.55e+05 3.11e+05 - 6.79e+06 (1.75e+05) -
1.46 $(16~{}separators)$ 2.99e+05 3.41e+05 - 7.66e+06 (2.32e+05) -
1.76 $(64~{}separators)$ 6.06e+05 5.98e+05 - 8.41e+06 (6.76e+04) -
vm1084 1.25 $(\mathcal{L}_{1}norm)$ 3.23e+05 3.61e+05 - 9.68e+06 (2.31e+05) -
1.21 $(4~{}separators)$ 2.70e+05 3.09e+05 - 8.89e+06 (1.89e+05) -
1.48 $(16~{}separators)$ 3.27e+05 3.80e+05 - 9.97e+06 (1.57e+05) -
1.73 $(64~{}separators)$ 7.30e+05 6.87e+05 - 1.06e+07 (2.37e+05) -
pcb1173 1.27 $(\mathcal{L}_{1}norm)$ 7.56e+04 8.36e+04 - 1.64e+06 (3.85e+04) -
1.20 $(4~{}separators)$ 6.41e+04 7.70e+04 - 1.49e+06 (2.59e+04) -
1.47 $(16~{}separators)$ 7.50e+04 8.84e+04 - 1.69e+06 (2.61e+04) -
1.67 $(64~{}separators)$ 1.52e+05 1.36e+05 - 1.79e+06 (1.72e+04) -
d1291 1.27 $(\mathcal{L}_{1}norm)$ 6.31e+04 6.85e+04 - 2.04e+06 (3.12e+04) -
1.23 $(4~{}separators)$ 6.15e+04 7.36e+04 - 1.88e+06 (2.83e+04) -
1.52 $(16~{}separators)$ 7.45e+04 8.50e+04 - 2.13e+06 (3.23e+04) -
1.75 $(64~{}separators)$ 1.60e+05 1.68e+05 - 2.25e+06 (1.61e+04) -
rl1304 1.27 $(\mathcal{L}_{1}norm)$ 3.32e+05 3.80e+05 - 1.10e+07 (2.56e+05) -
1.20 $(4~{}separators)$ 2.93e+05 3.35e+05 - 1.00e+07 (1.58e+05) -
1.48 $(16~{}separators)$ 3.57e+05 4.03e+05 - 1.13e+07 (1.24e+05) -
1.69 $(64~{}separators)$ 8.09e+05 7.36e+05 - 1.19e+07 (1.20e+05) -
rl1323 1.27 $(\mathcal{L}_{1}norm)$ 3.61e+05 4.20e+05 - 1.15e+07 (1.87e+05) -
1.22 $(4~{}separators)$ 3.17e+05 3.65e+05 - 1.05e+07 (1.48e+05) -
1.49 $(16~{}separators)$ 4.00e+05 4.23e+05 - 1.19e+07 (1.46e+05) -
1.72 $(64~{}separators)$ 8.16e+05 7.99e+05 - 1.26e+07 (1.36e+05) -
The first column of Table 2 lists the name of the instance, each formatted with an alphabetic prefix followed by a numeric value that indicates the number of points in that instance. The cause of its non-Euclidean characteristic is listed alongside the instance name, along with the resulting DF. It can be seen that the DF is positively correlated with the number of separators, though it is also dependent on the spatial distribution of points and their relative position with the added separators. The remaining columns list the solutions found using the various approaches, with the mean ($\mu$) and standard deviation ($\sigma$) listed for the metaheuristic solutions found at the end of one minute.
Figure 5: Ratio of heuristic solution costs to ACHCI cost
Figure 6: Computation time comparison
Figure 7: Ratio of metaheuristic solutions cost to ACHCI cost
To visualize the effect of DF, the ratio of tour costs obtained using the NN and NI heuristic to that obtained using the ACHCI algorithm is shown in Fig. LABEL:H_ratio_DF for each of the TSPLIB instances studied. It is clear that the ACHCI algorithm largely outperforms these competing heuristics, indicating that the proposed heuristic is well suited to a variety of non-Euclidean point cloud configurations. For increasing DF, a diminishing trend is observed in the advantage that the ACHCI algorithm provides, which is attributed to the increasing degree of approximation caused by the MDS when projecting points from an increasingly distorted non-Euclidean space to a Euclidean 2D space. It is also seen in Fig. LABEL:H_ratio_points that the relative performance of the ACHCI heuristic is largely unaffected by the number of points. The ACHCI tour cost was lower than NN and NI heuristic solutions in 88% and 99% of the cases studied respectively.
A worst case complexity of $O(n^{3})$ characterizes the ACHCI heuristic, as seen in Fig. 6 where the x axis is the number of points raised to $3$. This is attributed to the eigenvalue decomposition involved with MDS and the cheapest insertion criteria used when selecting points while building the tour. The NI heuristic also uses the cheapest insertion criteria but requires more computation time than the ACHCI heuristic because it is not initiated with the convex hull of points, resulting in an increased number of points for insertion. This results in no solution being found within the 60 second time limit for larger problem instances, as seen in Table 2. Because of the greedy nature of the NN heuristic, tours are produced almost instantaneously regardless of the number of points.
When comparing the performance of metaheuristic approaches, box plots enable the visualization of the stochasticity involved in the final solutions found. The relative performance of the GA and ACO solutions with respect to the ACHCI solutions decrease with increasing number of points as seen in Fig. LABEL:MH_ratio_points. This is because of the exponentially larger search space involved, and while the GA outputs solutions for any number of points because of the random generation of solutions, the ACO does not produce any solutions for larger instances. This is attributed to the resource intensive operation of constructing each ACO tour by sampling each edge from the initial transition probabilities. It can be seen however that the performance of GA and ACO for smaller problem sizes is comparable and the relative performance is mostly unaffected by the DF, as seen in Fig. LABEL:MH_ratio_DF. The ACHCI cost is lower than the mean GA and ACO costs in 87 % and 95 % of the studied test cases respectively.
6 Conclusion
The utilization of MDS enables the ACHCI heuristic to capitalize on the property that points on the boundary of the convex hull are visited in the same cyclic order as they appear in their optimal Euclidean TSP tour. While the initiating convex hull is based on an approximation of points in a Euclidean space, it is seen empirically that the non-Euclidean tours generated by the ACHCI heuristic outperform the benchmark algorithms. When compared with metaheuristic algorithms like the GA and ACO that iteratively update a population of solutions at high computational expense, the ACHCI tours outperformed them in a vast majority of cases, especially in larger test instances. ACHCI tour costs were lower than NI heuristic tour costs in 99% of the test cases studied, while also providing solutions more rapidly due to the reduced number of points for insertion. Though the NN heuristic is faster than the ACHCI heuristic because of its greedy approach to tour construction, the ACHCI tour costs were lower in 89% of the test cases. Considering the near instantaneous generation of NN tours, it may therefore be worthwhile to compute the NN tours in addition to the ACHCI tour, and simply choose the tour with the minimum cost.
Computational experiments presented in this paper were based on non-Euclidean test cases derived from TSPLIB instances by either the presence of obstacles, or the utilization of the $\mathcal{L}_{1}$ norm. While both these modifications are based on practical applications, it is easy to see that a similar performance can be expected from other real-world requirements that may arise out of other cost functions such as fuel consumption, emissions, travel time or driver comfort.
Acknowledgements
This work was supported by Ford Motor Company through the Ford-OSU Alliance Program. Declarations of interest: none